# DiffVLA++: Bridging Cognitive Reasoning and End-to-End Driving through Metric-Guided Alignment

### v2

@@ Abstract
Conventional end-to-end (E2E) driving models are effective at generating physically plausible trajectories, but often fail to generalize to long-tail scenarios due to the lack of essential world knowledge to understand and reason about surrounding environments. In contrast, Vision-Language-Action (VLA) models leverage world knowledge to handle challenging cases, but their limited 3D reasoning capability can lead to physically infeasible actions. In this work we introduce DiffVLA++, an enhanced autonomous driving framework that explicitly bridges cognitive reasoning and E2E planning through metric-guided alignment. First, we build a VLA module directly generating semantically grounded driving trajectories. Second, we design an E2E module with a dense trajectory vocabulary that ensures physical feasibility. Third, and most critically, we introduce a metric-guided trajectory scorer that guides and aligns the outputs of the VLA and E2E modules, thereby integrating their complementary strengths. The experiment on the ICCV 2025 Autonomous Grand Challenge leaderboard shows that DiffVLA++ achieves EPDMS of 49.12.
@@ 1 Introduction
This raises a key challenge: while E2E models excel at physically grounded trajectory prediction and VLA models excel at cognitive reasoning with world knowledge, a principled way to bridge these two paradigms has not been fully explored. The difficulty lies in combining semantically rich but sometimes physically infeasible VLA trajectories with physically plausible but semantically limited E2E trajectories.
@@
Evaluations on the ICCV 2025 Autonomous Grand Challenge benchmark show that DiffVLA++ achieves Extended Predictive Driver Model Score (EPDMS) [26] of 49.12.
@@ 2 Fully-Differentiable VLA Model for E2E Driving
The Vision-Language-Action (VLA) framework is designed around four core modules that jointly enable multimodal reasoning and trajectory generation.
The visual processing stream employs a CLIP-based Vision Transformer (ViT-L/14) that encodes multi-view images into a compact set of visual tokens. Each input image is resized and partitioned into patches, which are embedded into high-dimensional representations capturing spatial context. These tokens are then adapted through a Driving Vision Adapter that performs compression and projection, ensuring compatibility with the downstream language model.
@@
Finally, the last layer of the LLM projects the fused hidden states into continuous future trajectories of the ego vehicle. Instead of discretizing actions, the model directly predicts waypoints over a 4-second horizon, where each waypoint includes lateral position, longitudinal position, and heading angle. This fully differentiable design avoids discretization errors, improves smoothness, and allows end-to-end optimization of the driving policy.
@@ 3 Conventional End-to-End Driving Model
Following the multi-task architecture of [2], the E2E module comprises three prediction heads: (i) an agent detection head that regresses the states of dynamic agents, (ii) a semantic segmentation head that predicts scene semantics in BEV space, and (iii) a trajectory planning head that generates the ego vehicle’s future motion.
The agent detection head employs a set of learnable agent queries $\mathbf{Q}_{\text{agent}}\in\mathbb{R}^{N\times d}$, which interact with the BEV features via deformable cross-attention to produce agent-centric embeddings $\mathbf{f}_{a}\in\mathbb{R}^{N\times d}$. These embeddings are decoded into bounding box parameters $(x,y,w,l,\theta)$, representing the center coordinates, width, length, and heading of each detected agent. Together with the semantic segmentation output, this ensures that the BEV representation encodes both dynamic and contextual information for downstream planning (Sec. 3.1) and metric-guided alignment (Sec. 4).
@@ 3.1 Trajectory Planning Head
The trajectory planning head operates over a pre-defined dense trajectory vocabulary
$\mathcal{V}\subset\mathbb{R}^{8192\times 8\times 3},$
where each candidate trajectory $v\in\mathcal{V}$ consists of 88 waypoints sampled at $2\,\text{Hz}$ over a four-second horizon. Each waypoint $\mathbf{p}_{t}=(x_{t},y_{t},\theta_{t})\in\mathbb{R}^{3}$ encodes the 2D position and heading angle at time step $t$. The vocabulary $\mathcal{V}$ is constructed via $K$-means clustering on expert trajectories from the navtrain split from Navsim dataset, ensuring broad coverage of feasible motion patterns.
For each candidate trajectory $v=\{\mathbf{p}_{t}\}_{t=0}^{7}$, BEV features are sampled at the waypoint locations using bilinear grid sampling. These features are aggregated into a compact trajectory-level embedding $\mathbf{f}_{v}\in\mathbb{R}^{d}$ via a learned fusion mechanism:
$\begin{gathered}\mathbf{z}_{v}=\big[\mathbf{F}(\mathbf{p}_{0}),\dots,\mathbf{F}(\mathbf{p}_{7})\big]\in\mathbb{R}^{8\times d},\\ w=\mathrm{MLP}(\mathbf{s}_{\text{ego}})\in\mathbb{R}^{8},\\ \mathbf{f}_{v}=\sum_{t=0}^{7}w_{t}\,\mathbf{z}_{v}^{(t)},\end{gathered}$
where $\mathbf{F}(\mathbf{p}_{t})$ denotes the BEV feature vector at position $(x_{t},y_{t})$, $\mathbf{z}_{v}$ is the sequence of sampled waypoint features, $\mathbf{s}_{\text{ego}}$ represents the current ego state (e.g., speed, acceleration, and high-level command concatenated and encoded using a MLP), and $w_{t}$ is attention weight for t-th waypoint derived from $\mathbf{s}_{\text{ego}}$ via an MLP. This design allows the fusion mechanism to emphasize waypoints that are more relevant to the current driving context.
To incorporate dynamic scene context, the trajectory embedding $\mathbf{f}_{v}$ is further refined through cross-attention with both agent features $\mathbf{f}_{a}$ and the ego state $\mathbf{s}_{\text{ego}}$:
$\hat{\mathbf{f}}_{v}=\texttt{CrossAttn}\!\left(\texttt{CrossAttn}(\mathbf{f}_{v},\mathbf{f}_{a}),\mathbf{s}_{\text{ego}}\right),$
where $\hat{\mathbf{f}}_{v}$ denotes the context-enriched trajectory representation.
Finally, a decoding MLP transforms $\hat{\mathbf{f}}_{v}$ into a residual offset vector $\Delta v\in\mathbb{R}^{8\times 3}$, comprising per-waypoint corrections $(\Delta x_{t},\Delta y_{t},\Delta\theta_{t})$. The final output trajectory is obtained by:
$v_{\text{pred}}=v+\Delta v.$
In addition, the trajectory embedding $\mathbf{f}_{v}$ is also used by the metric-guided trajectory scorer introduced in Sec. 4.
@@ 4 Metric-Guided Alignment
A key novelty of DiffVLA++ lies in metric-guided alignment, which serves as the bridge between the cognitively rich but sometimes physically infeasible trajectories from the VLA module and the physically grounded but semantically limited outputs of the E2E module. To achieve this, we employ a lightweight trajectory scorer that maps trajectory features into explicit driving metrics, thereby providing a shared evaluation space for both systems.
The trajectory scorer is implemented as a set of parallel MLP heads, each regressing one driving metric from the Navsim simulator. Given a trajectory feature $\mathbf{f}_{v}$, the scorer predicts $\hat{s}_{m}$ for metric $m$:
$\hat{s}_{m}=\mathrm{MLP}_{m}(\mathbf{f}_{v}),\quad m\in\begin{cases}\text{NC},\text{DAC},\text{DDC},\text{TLC},\\ \text{EP},\text{TTC},\text{LK},\text{HC}.\end{cases}$ (1)
The simulator provides eight driving performance metrics, categorized as follows:
  • EP (Ego Progress): continuous score in $[0,1]$ measuring progress along the route centerline.
  • DAC (Drivable Area Compliance), TLC (Traffic Light Compliance), TTC (Time-to-Collision), LK (Lane Keeping), HC (Hazard Compliance): binary scores in $\{0,1\}$ indicating whether the ego vehicle violates specific driving rules.
  • NC (Navigation Compliance), DDC (Dynamic Object Distance Compliance): ternary scores in $\{0,0.5,1\}$, where intermediate penalties are assigned when infractions are not directly caused by the ego vehicle.
The scorer is trained jointly with the E2E driving model, thereby correlating the BEV feature space with rule-based driving evaluations and providing auxiliary supervision for safer trajectory selection. Ground-truth metric labels are collected from the navtrain split of the Navsim dataset. Training minimizes a weighted regression objective:
$\mathcal{L}_{\hat{s}}=\sum_{i,m}w_{m}\left\|\hat{s}_{m}^{i}-s_{m}^{i}\right\|_{2}^{2},$
where $\hat{s}_{m}^{i}$ and $s_{m}^{i}$ denote the predicted and ground-truth scores for metric $m$ and the $i$-th trajectory, respectively, and $w_{m}$ are per-metric weights balancing scale and importance.
By projecting both VLA-generated and E2E-generated trajectories into the same metric space, this scorer enables explicit alignment of VLA trajectories and E2E trajectories.
@@ 5 Post-Processing
$s_{\text{final}}^{\text{E2E}}$ $=w_{1}\cdot\hat{s}_{\text{NC}}+w_{2}\cdot\hat{s}_{\text{DAC}}+w_{3}\cdot\hat{s}_{\text{EP}}$ (2)
@@ 6.1 Training for VLA
We train the VLA module end-to-end and the model directly regresses future waypoints over a 4-second horizon at 2 Hz, where each waypoint includes $(x,y,\theta)$. Training is performed using AdamW [30] with a cosine learning rate schedule [31], an initial learning rate of $1\times 10^{-5}$. A dropout rate of 0.05 is applied to both the vision adapter and the LLM. The VLA model is trained for 1 epoch with a batch size of 8 across 8 NVIDIA A800 GPUs.
@@ 6.2 Training for E2E and scorer in Metric-Guided Alignment
The E2E module and the trajectory scorer are trained jointly to ensure a consistent BEV feature space that captures both semantic and dynamic scene information. The overall training objective includes the following components: agent bounding box regression loss, agent classification loss, semantic segmentation loss, trajectory imitation loss, and scorer loss. We assign loss weights as follows: $1.0$ for agent bounding box regression, $10.0$ for agent classification, $20.0$ for trajectory imitation, $14.0$ for semantic segmentation, and $14.0$ for the scorer. The model is trained for $30$ epochs with a total batch size of 88 and an initial learning rate of $1\times 10^{-4}$ on 8 A800 GPUs. The E2E module use same optimizer and learning rate schdular as VLA module.

### v3

@@ Abstract
Conventional end-to-end (E2E) driving models are effective at generating physically plausible trajectories, but often fail to generalize to long-tail scenarios due to the lack of essential world knowledge to understand and reason about surrounding environments. In contrast, Vision-Language-Action (VLA) models leverage world knowledge to handle challenging cases, but their limited 3D reasoning capability can lead to physically infeasible actions. In this work we introduce DiffVLA++, an enhanced autonomous driving framework that explicitly bridges cognitive reasoning and E2E planning through metric-guided alignment. First, we build a VLA module directly generating semantically grounded driving trajectories. Second, we design an E2E module with a dense trajectory vocabulary that ensures physical feasibility. Third, and most critically, we introduce a metric-guided trajectory scorer that guides and aligns the outputs of the VLA and E2E modules, thereby integrating their complementary strengths. The experiment on the ICCV 2025 Autonomous Grand Challenge leaderboard shows that our model achieves EPDMS of 49.12.
@@ 1 Introduction
This raises a key challenge: while E2E models excel at grounded trajectory prediction and VLA models excel at cognitive reasoning with world knowledge, a principled way to bridge these two paradigms has not been fully explored. The difficulty lies in combining semantically rich yet occasionally infeasible VLA trajectories with physically plausible but semantically limited E2E trajectories.
@@
Evaluations on the ICCV 2025 Autonomous Grand Challenge benchmark show DiffVLA++ achieves Extended Predictive Driver Model Score (EPDMS) [26] of 49.12.
@@ 2 Fully-Differentiable VLA Model for E2E Driving
The Vision-Language-Action (VLA) framework is designed around four core modules that jointly enable multimodal reasoning and trajectory generation. The visual processing stream employs a CLIP-based Vision Transformer (ViT-L/14) that encodes multi-view images into a compact set of visual tokens. Each input image is resized and partitioned into patches, which are embedded into high-dimensional representations capturing spatial context. These tokens are then adapted through a Driving Vision Adapter that performs compression and projection, ensuring compatibility with the downstream language model.
@@
Finally, the last layer of the LLM projects the fused hidden states into continuous future trajectories of the ego vehicle. Instead of discretizing actions, the model directly predicts waypoints over a four second horizon, where each waypoint includes lateral position, longitudinal position, and heading angle. This fully differentiable design avoids discretization errors, improves smoothness, and allows end-to-end optimization of the driving policy.
@@ 3 Conventional End-to-End Driving Model
Following the multi-task architecture of Transfuser [2], the E2E module comprises three prediction heads: (i) an agent detection head that regresses the states of dynamic agents, (ii) a semantic segmentation head that predicts scene semantics in BEV space, and (iii) a trajectory planning head that generates the ego vehicle’s future motion.
The agent detection head employs a set of learnable agent queries $\mathbf{Q}_{\text{agent}}\in\mathbb{R}^{N\times d}$ with $N=32$. Each query interacts with the BEV features via deformable cross-attention to produce agent-centric embeddings $\mathbf{F}_{a}\in\mathbb{R}^{N\times d}$. These embeddings are decoded into bounding box parameters $(x,y,w,h,\theta)$, representing the center coordinates, width, length, and heading of each detected agent. Together with the semantic segmentation output, this ensures that the BEV representation encodes both dynamic and contextual information for downstream planning (Sec. 3.1) and metric-guided alignment (Sec. 4).
@@ 3.1 Trajectory Planning Head
The trajectory planning head operates over a pre-defined dense trajectory vocabulary $\mathcal{V}=\{v_{i}\}_{i=1}^{M}$ with $M=8192$, where each candidate $v_{i}$ consists of eight waypoints sampled at $2\,\text{Hz}$ over a four second horizon. Each waypoint $\mathbf{p}_{t}=(x_{t},y_{t},\theta_{t})\in\mathbb{R}^{3}$ encodes the 2D position and heading angle at time $t$. The vocabulary is constructed via $K$-means clustering on expert trajectories from the navtrain split of the Navsim dataset, ensuring comprehensive coverage of feasible motion patterns.
To encode each candidate trajectory, BEV features are sampled at its waypoints using bilinear grid sampling. This yields a feature sequence for each $v_{i}$, which is aggregated into a compact embedding through a learned attention mechanism conditioned on the ego state. Collectively, these operations produce a set of trajectory embeddings $\mathbf{F}_{v}=[\mathbf{f}_{1},\dots,\mathbf{f}_{M}]^{\top}\in\mathbb{R}^{M\times d}$, where each $\mathbf{f}_{i}\in\mathbb{R}^{d}$ summarizes the visual context along $v_{i}$.
To incorporate dynamic scene context, $\mathbf{F}_{v}$ is refined by attending to the agent-centric features $\mathbf{F}_{a}\in\mathbb{R}^{N\times d}$ from the detection head:
$\mathbf{F}_{v}^{\text{ctx}}=\texttt{CrossAttn}(\mathbf{F}_{v},\mathbf{F}_{a},\mathbf{F}_{a}),$
where CrossAttn denotes deformable cross-attention, enabling each trajectory query to attend to relevant agent features. The resulting $\mathbf{F}_{v}^{\text{ctx}}\in\mathbb{R}^{M\times d}$ represents a context-aware set of trajectory hypotheses.
Each refined embedding $\mathbf{f}_{i}^{\text{ctx}}\in\mathbf{F}_{v}^{\text{ctx}}$ is then decoded into a residual offset $\Delta v_{i}\in\mathbb{R}^{8\times 3}$ via an MLP, yielding the final predicted trajectory:
$v_{\text{pred}}=v_{i}+\Delta v_{i},$
with $v_{i}$ selected based on downstream scoring (Sec. 4). The context-aware embeddings $\mathbf{F}_{v}^{\text{ctx}}$ are also shared with the metric-guided scorer for joint optimization. $d$ is set to 256 for all modules.
@@ 4 Metric-Guided Alignment
A key novelty of DiffVLA++ lies in metric-guided alignment, which serves as the bridge between the cognitively rich yet occasionally physically inconsistent trajectories from the VLA module and the physically grounded but semantically limited outputs of the E2E module. To achieve this, we employ a lightweight trajectory scorer that maps trajectory features into explicit driving metrics, thereby providing a shared evaluation space for both systems.
The trajectory scorer is implemented as a set of parallel MLP heads, each regressing one driving metric from the Navsim simulator. Given the context-aware trajectory embeddings $\mathbf{F}_{v}^{\text{ctx}}=[\mathbf{f}_{1}^{\text{ctx}},\dots,\mathbf{f}_{M}^{\text{ctx}}]^{\top}\in\mathbb{R}^{M\times d}$ produced by the planning head (Sec. 3.1), the scorer simultaneously predicts metric scores $\hat{s}_{m}^{i}$ for each candidate $v_{i}$ and metric $m$:
$\hat{s}_{m}^{i}=\mathrm{MLP}_{m}(\mathbf{f}_{i}^{\text{ctx}}),\quad m\in\begin{subarray}{c}\{\text{NC},\text{DAC},\text{DDC},\text{TLC},\\ \text{EP},\text{TTC},\text{LK},\text{HC}\}\end{subarray}.$
Here, the eight driving metrics are categorized as follows:
  • EP: continuous score in $[0,1]$ measuring progress along the route centerline.
  • DAC, TLC, TTC, LK, HC: binary scores in $\{0,1\}$ indicating compliance with drivable area, traffic lights, collision avoidance, lane keeping, and history comfort.
  • NC, DDC: ternary scores in $\{0,0.5,1\}$ measuring collision and driving direction, where intermediate penalties are assigned when infractions are not directly caused by the ego vehicle.
The scorer is trained jointly with the E2E driving model, correlating the BEV feature space with rule-based driving evaluations and providing auxiliary supervision for safer trajectory selection. Ground-truth labels are collected from the navtrain split of the Navsim dataset. Training minimizes a weighted composite objective:
$\mathcal{L}_{\hat{s}}=\sum_{i=1}^{M}\sum_{m}w_{m}\,\ell_{m}(\hat{s}_{m}^{i},s_{m}^{i}),$
where $\hat{s}_{m}^{i}$ and $s_{m}^{i}$ denote the predicted and ground-truth scores for metric $m$ and candidate $v_{i}$, respectively, $w_{m}$ are per-metric weights balancing scale and importance, and $\ell_{m}$ is the task-specific loss function:
  • MSE for continuous metrics (EP);
  • Binary Cross-Entropy (BCE) for binary metrics (DAC, TLC, TTC, LK, HC);
  • Cross-Entropy for ternary metrics (NC, DDC).
By projecting both VLA-generated and E2E-generated trajectories into this shared metric space, the scorer enables explicit alignment through a common, interpretable performance benchmark.
@@ 5 Post-Processing
$s_{\text{final}}^{\text{E2E}}$ $=w_{1}\cdot\hat{s}_{\text{NC}}+w_{2}\cdot\hat{s}_{\text{DAC}}+w_{3}\cdot\hat{s}_{\text{EP}}$
@@ 6.1 Training for VLA
We train the VLA module end-to-end and the model directly regresses future waypoints over a 4-second horizon at 2 Hz, where each waypoint includes $(x,y,\theta)$. Training is performed using AdamW [30] with a cosine learning rate schedule [31], an initial learning rate of $1\times 10^{-5}$. A dropout rate of 0.05 is applied to both the vision adapter and the LLM. The VLA model is trained for one epoch with a batch size of $8$ across eight NVIDIA A800 GPUs.
@@ 6.2 Training for E2E and scorer in Metric-Guided Alignment
The E2E module and the trajectory scorer are trained jointly to ensure a consistent BEV feature space that captures both semantic and dynamic scene information. The overall training objective includes the following components: agent bounding box regression loss, agent classification loss, semantic segmentation loss, trajectory imitation loss, and scorer loss. We assign loss weights as follows: $1.0$ for agent bounding box regression, $10.0$ for agent classification, $20.0$ for trajectory imitation, $14.0$ for semantic segmentation, and $14.0$ for the scorer. The model is trained for $30$ epochs with a total batch size of $8$ and an initial learning rate of $1\times 10^{-4}$ on four A800 GPUs. The E2E module use same optimizer and learning rate schdular as VLA module.